Mattola Company is giving each of its employees a holiday bonus of $100 on December 13 (a nonpayday). The company wants each employee's check to be $100. The supplemental tax percent is used. Nobody has capped for OASDI prior to the bonus check. a. What will be the gross amount of each bonus if each employee pays a state income tax of 2.8% (besides the other payroll taxes)? You may need to add one penny to the gross so that net bonus is exactly $100.00. Round your calculations and final answers to the nearest cent. $ 154.93

b. What would the net amount of each bonus check be if the company did not gross-up the bonus? $ 64.56

These answers are wrong and i do not see what im doing wrong
Answer #1

Hi, Friend...

Hope you are also doing well.

I have tried to give the best possible answer. If you have any doubts, please mention in the comment section. I'll definitely respond in early possible time.

Ans:

a ) Calculation of gross amount of each bonus:

state income tax = 2.8 %

FICA social security tax = 6.20 %

FICA Medicare tax = 1.45 %

gross amount of each bonus = $100/ ( 100 % - 25 % - 6.20 % - 1.45 % - 2.8 % )

= 100 / 64.55 %

= $154.9187 or we can say $154.92

b ) Calculation of net amount of each bonus:

= 100 * ( 100 % - 25 % - 6.20 % - 1.45 % - 2.80 % )

= 100 * 64.55 %

= $64.55

Notes:

Supplemental tax percent is: 25%

FICA-social security tax is: 6.2%

FICA-Medicare tax is: 1.45%.

State income tax is: 2.8%

Thanks, Dear... Have a Wonderful Day.